OpenStudy (anonymous):

m/4+7=-7 two step equation show work

OpenStudy (anonymous):

m/4 = -14 m = - 56

OpenStudy (annon):

m/4=-7-7 m/4=-14 m=-14*4=-56

OpenStudy (anonymous):

Subtract 7 from both sides m/4 = -14 multiply both sides by 4 m = -56